TheSmartOne (thesmartone):

Calculus 1 Question Find the equations of both tangent lines to the graph of the ellipse x^2/4 + y^2/9 = 1 that passes through the point (0, 4) not on the graph.

TheSmartOne (thesmartone):

\(\Large \frac{x^2}{4} + \frac{y^2}{9} = 1\)

TheSmartOne (thesmartone):

That equation is the same thing as 9x^2 + 4y^2 = 36 And differentiating that, we get 18x + 8yy' = 0 Solving for y' we get y' = (-9/4) * (x/y) We know the slope is the same as the derivative (y - 4)/(x - 0) = (-9/4) * (x/y) (y-4)/x = (-9x)/(4y) Cross multiply 4y^2 - 16y = -9x^2 Same thing as 9x^2 + 4y^2 = 16y And plugging in what we know from the beginning 36 = 16y y = 2.25 And I'm not sure what to do after this >.>

TheSmartOne (thesmartone):

Maybe plug it back in to the ellipse? Lemme try plugging that back in to the ellipse 9x^2 + 4y^2 = 36 9x^2 + 4(2.25)^2 = 36 9x^2 + 20.25 = 36 x^2 = (36 - 20.25)/9 x^2 = 1.75 \(\Large x = \pm \sqrt{1.75}\)

TheSmartOne (thesmartone):

If you look at this graph https://www.desmos.com/calculator/se3bojn9xg How would a tangent line go through the point (0, 4) and have a slope of 0?

OpenStudy (visarika23):

Ya I was thinking the same

OpenStudy (ajspeller):

Yeah I just graphed it myself. I think the problem has an error in it.

OpenStudy (ajspeller):

The point should be on the ellipse.

TheSmartOne (thesmartone):

 We need to find those two tangent lines, I believe

OpenStudy (ajspeller):

oh okay. Sorry I mis-read the question

OpenStudy (s4sensitiveandshy):

when x=-1.75, 1.75 y = ???

TheSmartOne (thesmartone):

It's srt(1.75) and -sqrt(1.75) and y would equal 2.25 or -2.25

TheSmartOne (thesmartone):

but we would want the positive y-value so 2.25

OpenStudy (visarika23):

Is the eq of tangent y=+/-1.322x+4 and the point of contact is (1.32,2.25) and (-1.32,2.25)

TheSmartOne (thesmartone):

that's what I got, thanks c:

OpenStudy (visarika23):

☺☺